OpenStudy (anonymous):

Complete the equation for the linear function whose graph contains the points (−1, −2) and (3, 10). y + _ = _ (x + 1)

pooja195 (pooja195):

\(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\sf\color{blue}{~~~~Welcome~to~OpenStudy!~\ddot\smile}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

pooja195 (pooja195):

\[y1-y2=m(x-x1)\] plug in the numbers :)

pooja195 (pooja195):

Tell me what you get :)

pooja195 (pooja195):

x1 y1 x2 y2 \[(−1, −2) and (3, 10). \]

OpenStudy (anonymous):

I'm sorry I don't understand.

pooja195 (pooja195):

You have the point slope formula now u just have to plug in the X and Y values

OpenStudy (anonymous):

ok so x -1 and x3, y -2 and y10?

pooja195 (pooja195):

yes so it would look like \[y -2 = m (x + 1) \]

OpenStudy (anonymous):

Thanks so much!